OpenStudy (anonymous):

Can anyone show me how to prove the following trig identity (sin(3x)/(sinx)-(cos(3x)/cosx)=2

OpenStudy (anonymous):

Is this complete?? I mean where is LHS and where is RHS.. You want to prove this equal to what thing??

OpenStudy (anonymous):

equals 2 I am sorry I forgot to post that.....

OpenStudy (anonymous):

See: Take the LCM first and combine the terms: \[\frac{\sin(3x)}{\sin(x)} - \frac{\cos(3x)}{\cos(x)} \implies \frac{\sin(3x) \cdot \cos(x) - \cos(3x) \cdot \sin(x)}{\sin(x)\cos(x)}\] \[\implies \frac{\sin(3x - x)}{\sin(x) \cdot \cos(x)} \implies \frac{\sin(2x)}{\sin(x) \cdot \cos(x)} \implies \frac{2\sin(x) \cdot \cos(x)}{\sin(x) \cdot \cos(x)} \implies 2\]

OpenStudy (anonymous):

Oh I see now thanks you for your response.

OpenStudy (anonymous):

Identities used are: \[\sin(A) \cdot \cos(B) - \cos(A) \cdot \sin(B) = \sin(A-B)\] \[\sin(2A) = 2\sin(A) \cdot \cos(A)\]

OpenStudy (anonymous):

Welcome dear..